3x + 17 - 5x = 12 - (6x + 3)
Answer by checkley71(8403)

3X+17-5X=12-(6X+3)
3X-5X+17=12-6X-3
-2X+6X=-17+12-3
4X=-8
X=-8/4
X=-2
PROOF
3*-2+17-5*-2=12-(6*-2+3)
-6+17+10=12-(-12+3)
21=12+12-3
21=21